why is the equivalent weight of mohr’s salt 1?

Grade:12

4 Answers

ankit goyal
8 years ago
equi. wt. of mohr salt = 392  because
Equivalent weight of Mohr salt = Mol. Wt./ no. of equivalents. = 392/ 1. = 392 .
approve if useful
Aarti Gupta
8 years ago
Mohr’s salt is ferrous ammonium sulphate and its an reducing agent.The equivalent weight of Mohr’s salt is 392 which is equal to its molecular weight.This is due to the fact that oxidation of ferrous(Fe2+) to ferric(Fe3+) involves only one electron change only.
Fe2+ ----------> Fe3+ + e--
Equivalent weight = Molecular weight/1 = 392/1 = 392
Sai Sampath
8 years ago
Mohr’s salt is ferrous ammonium sulphate and its an reducing agent.The equivalent weight of Mohr’s salt is 392 which is equal to its molecular weight equi. wt. of mohr salt = 392  becauseEquivalent weight of Mohr salt = Mol. Wt./ no. of equivalents. =392/ 1. = 392 .
 
Nishanth
4 years ago
As Mohr salt is a double salt. In aqueous solution Mohr salt gives ammonium ion ferrous ion and sulphate ions.It can act only as a reducing agent i.e., ferrous to ferric ion . So n factor is 1 . Equivalent weight is 392